OpenStudy (anonymous):

The weight of 1000 identical samples of a substance is 10 pounds. What is the weight of 10 samples?

OpenStudy (anonymous):

is it 100??? idk. lol.

OpenStudy (anonymous):

wait, would it be 0.1?

OpenStudy (lgbasallote):

hmm it depends...do you mean the combined weight of 1000 identical samples?

OpenStudy (anonymous):

No clue, that's the exact question/problem lol.

OpenStudy (lgbasallote):

it smells like a trick question...it depends on the interpretation.. the weight of 1000 identical samples may mean that the weight of 1000 samples are all 10 lbs...or that the combined weight of these is 10 lbs...so im going to give you 2 equations...one is that all of them weigh 10 lbs...the other would be... 1000x = 10 so x = 10/1000 or 1/100 so 10 of these would be 10(1/100) = 1/10 = 0.1

OpenStudy (anonymous):

I'll try 0.1 and cross my fingers! haha. (It's an online course)

OpenStudy (lgbasallote):

lol cross it well :P because the way i interpreted this is that it's a trick question